OpenStudy (anonymous):

Convert the following points in polar coordinates to rectangular coordinates: a.(1, 240°) b.(4, 135°) c.(–3, 285°) d.(10, 20°) Convert the following rectangular equations to polar form: a. y = 5 b. (x – 1)2 + y2 = 1 c. 5x + 4y = 2

OpenStudy (johnweldon1993):

Hint* you are given coordinates in the form \(\large (r, \theta)\) \[\large x = rcos(\theta)\] \[\large y = rsin(\theta)\]